A plumber needs four sections of pipe 2 7/8 feet long. Can this be cut from a 12-foot section? If so, how much pipe will be left over? If not, why not?
Answer by stanbon(75887)

A plumber needs four sections of pipe 2 7/8 feet long. Can this be cut from a 12-foot section? If so, how much pipe will be left over? If not, why not?
-----------------
12 ft - 4(2 7/8) ft.
= 12 - 4(23/8)
= 12 - 92/8
= 96/8 - 92/8 = 4/8= 1/2 ft.
--------------------
There will be 1/2 ft left over.
==========================
Cheers,
Stan H.